Hi, I need help to solve 3x – 5y = 15 by the substitution method.
y = 2x + 4

Answer by stanbon(57329)

solve 3x – 5y = 15 by the substitution method.
y = 2x + 4
----
Substitute for "y" and solve for "x":
3x-5(2x+4) = 15
3x-10x-20 = 15
-7x = 35
x = -5
----
Solve for "y":
y = 2x+4
y = 2*-5+4
y = -6
=============
Cheers,
Stan H.
=============